2. Order: penicillin G potassium 125,000 units IV q6h. The infant weighs
2,500 grams. The recommended dose for infants is 150,000–300,000
units/kg/d divided in equal doses q4–6h. Is the ordered dose safe?
Answer #1

Recommended dose = 150000-300000units/Kg/day

child's weight = 2500grams = 2.5Kg

So for the child max recommended dose per day is 300000 × 2.5

                                                                    = 750000 Units

If given 6 hourly, one dose will be = 750000/4 = 1,87,000 Units

The dose ordered = 1,25000 Units

The ordered dose is well within the limits. So the dose is safe.